I'm supposed to solve these systems of equations. they are:
-5x + 4y= 3
x= 2y - 15
I'm supposed to find:
X=
y=
please help me. iv'e been trying to figure this out for a while. I don't know where to start because the second equation is throwing me off.

Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

solve these systems of equations. they are:
-5x + 4y= 3
x= 2y - 15
-----
Use the 2nd equation.
Substitute for "x" in the 1st equation to get:
-5(2y-15) + 4y = 3
-10y+75 + 4y = 3
-6y = -72
y = 12
----
Now solve for "x"::
x = 2y-15
x = 2*12-15
x = 24-15
x = 9
------------
Cheers,
Stan H.
-----------

Answer by ikleyn(52786)

.
On solving systems of two equations in two unknowns see the lessons
    - Solution of a linear system of two equations in two unknowns by the Substitution method
    - Solution of a linear system of two equations in two unknowns by the Elimination method
    - Solution of a linear system of two equations in two unknowns using determinant
    - Geometric interpretation of a linear system of two equations in two unknowns
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Systems of two linear equations in two unknowns".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.